OpenStudy (babydoll.m):

The length of the hypotenuse of a right triangle is 24 inches. If the length of one leg is 8 inches, what is the approximate length of the other leg? 16.0 inches 18.4 inches 22.6 inches 25.3 inches

OpenStudy (anonymous):

since it is a right triangle, you can use the pythagorean theorem.

OpenStudy (babydoll.m):

how do I use that

OpenStudy (anonymous):

i suggest drawing a picture, labelling it with the info, writing an appropriate equation and then solving for what you need.

OpenStudy (babydoll.m):

k

OpenStudy (anonymous):

 \[a^2+b^2 = c^2\]

OpenStudy (babydoll.m):

still don't understand

OpenStudy (babydoll.m):

25.3? @pgpilot326

OpenStudy (anonymous):

 \[\Rightarrow b^2 = c^2 - a^2 \Rightarrow b^2 = 24^2 - 8^2 = 576-64=512\]so \(b= \sqrt{512}\), right?

OpenStudy (anonymous):

remember, the hypotenuse is always opposite the right angle

OpenStudy (babydoll.m):

thank you so much

OpenStudy (anonymous):

you're welcome